# Tight Bounds for Cycle-Edge Decompositions and Covers

### v1

@@ Abstract.
An old conjecture of Gallai states that every $n$ vertex graph can be decomposed, that is $E(G)$ can be partitioned, into $\lceil\frac{n}{2}\rceil$ or fewer paths. In response to this conjecture, Lovász proved in 1968 that every $n$ vertex graph can be decomposed into $\lfloor\frac{n}{2}\rfloor$ or fewer cycles and paths. We study a related conjecture which states that every $n$ vertex graph can be decomposed into n1n-1 or fewer cycles and edges. We prove this conjecture for all graphs with maximum degree at most $4$. We also show that every $n$ vertex claw-free graph can be decomposed into n1n-1 or fewer $2$-regular subgraphs and edges. In addition, we consider the analogues covering problem. We prove that every graph $G$ containing a cycle can be covered by n2n-2 or fewer cycles and edges.
@@ 1. Introduction
The primary focus on this paper will be a related conjecture of Beikmohammadi, which is unpublished prior to this paper.
Conjecture 1.2.
Every $n$ vertex graph $G$ can be decomposed into at most n1n-1 cycles and edges.
This conjecture should be viewed as a version of Lovász’ Theorem for decompositions into cycles and edges rather than cycles and paths. Notice that if correct, this bound is tight for every tree. We also consider several relaxations of this conjecture.
To this end, we define the following parameters. For a graph $G$, let $\operatorname{f_{ce}}(G)$ denote the minimum number of subgraphs that needed to decompose $G$ into cycles and edges, and let $\operatorname{f_{re}}(G)$ denote the minimum number of subgraphs that is needed to decompose $G$ into $2$-regular subgraphs and edges. Similarly, let $g_{ce}(G)$ be the minimum number of subgraphs needed to cover $G$ by cycles and edges.
Given these notations, we can restate Conjecture 1.2 as asking if for all $n$ vertex graphs $G$, $\operatorname{f_{ce}}(G)\leq n-1$. In the same vein as Bonamy and Perrett’s result for Gallai’s Conjecture on graphs with maximum degree $5$, we prove Conjecture 1.2 holds for graphs with maximum degree at most $4$.
@@
Next, we prove that the analogue of Conjecture 1.2 for decompositions into $2$-regular graphs and edges, is true for claw-free graphs. Notice this bound is still tight for all $n$, as demonstrated by paths. Here if $G$ is $H$-free, then $G$ does not have $H$ as an induced subgraph. Meanwhile, as is standard, the claw denotes the graph $K_{1,3}$.
@@
The final set of relaxations of Conjecture 1.2 we consider is to study covers by cycles and edges, rather than decompositions into cycles and edges. Note the best possible upper bound for all graphs is still n1n-1 due to trees. We are able to show the only graphs to require n1n-1 cycles or edges in a cover are trees.
@@ 6. Future Work
Show Conjecture 1.2 holds if $G$ has maximum degree $5$, or if $G$ is $6$-regular, or $8$-regular.
@@
For all $n$ vertex graphs $G$, $\operatorname{f_{re}}(G)\leq n-1$.
@@ Acknowledgement
The research visit of S. Akbari at Simon Fraser University was supported in part by the ERC Synergy grant (European Union, ERC, KARST, project number 101071836). Clow is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) through PGS D-601066-2025.

### v2

@@ Abstract.
An old conjecture of Erdős and Gallai states that every $n$ vertex graph can be decomposed, that is $E(G)$ can be partitioned, into $O(n)$ cycles and edges. The covering version of this conjecture was proven by Pyber in 1985, where it was shown that all graphs can be covered by n1n-1 cycles and edges. The best upper bound on the number of cycles and edges required to decompose any graph is $O(n\log^{*}(n))$, which was recently shown by Bucić and Montgomery in 2023. Here $\log^{*}(n)$ denotes the iterated logarithm function. Meanwhile, a construction of Erdős demonstrate that there exists graphs which require $(\frac{3}{2}-o(1))n$ cycles and edges to be decomposed. We prove all graphs with maximum degree at most $4$ can be decomposed into n1n-1 or fewer cycles and edges. We also show that every $n$ vertex claw-free graph can be decomposed into n1n-1 or fewer $2$-regular subgraphs and edges. Finally, we prove that every graph $G$ containing a cycle can be covered by n2n-2 or fewer cycles and edges. This improves Pyber’s covering theorem by proving that n1n-1 cycles and edges are required only for trees.
@@ 1. Introduction
The primary focus on this paper will be a related conjecture of Erdős and Gallai. This conjecture appears in multiple of Erdős’ problem collections [9, 10, 11, 12] and has also been highlighted by other authors [3, 6, 7, 18, 19, 21, 25, 26].
Conjecture 1.2 (The Erdős-Gallai Conjecture [13]).
Every $n$ vertex graph $G$ can be decomposed into $O(n)$ cycles and edges.
Constructions, first by Gallai [13] and later by Erdős [12] demonstrate that $(1+o(1))n$ cycles and edges are insufficient for general graphs. In particular, Erdős [12] demonstrates a graph where $(\frac{3}{2}-o(1))n$ cycles and edges are required in a decomposition.
A simple argument, observed by Erdős and Gallai in the 1960s, consisting of iteratively removing longest cycles gives an upper bound of $O(n\log(n))$. The first improvement on this bound came in 2014 from Conlon, Fox, and Sudakov [7] who showed $O(n\log\log(n))$ cycles and edges are sufficient. This was recently improved to $O(n\log^{*}(n))$ in 2023 by Bucić and Montgomery [6]. Here $\log^{*}(n)$ denotes the iterated logarithm of $n$. The only cases where the Erdős-Gallai conjecture has been proven is typical binomial random graphs [7, 19, 21], and graphs with linear minimum degree [7, 18].
We are especially interested in graph classes where exactly n1n-1 or fewer cycles and edges are sufficient. To this end, we define the following parameters. For a graph $G$, let $\operatorname{f_{ce}}(G)$ denote the minimum number of subgraphs that needed to decompose $G$ into cycles and edges, and let $\operatorname{f_{re}}(G)$ denote the minimum number of subgraphs that is needed to decompose $G$ into $2$-regular subgraphs and edges. Similarly, let $g_{ce}(G)$ be the minimum number of subgraphs needed to cover $G$ by cycles and edges.
Given these notations, we can restate Conjecture 1.2 as asking if for all $n$ vertex graphs $G$, $\operatorname{f_{ce}}(G)=O(n)$. In the same vein as Bonamy and Perrett’s result for Gallai’s Conjecture on graphs with maximum degree $5$, we prove that all graphs with maximum degree at most $4$ can be decomposed with n1n-1 cycles and edges.
@@
Next, we prove that all claw-free graphs can be decomposed into at most n1n-1, 22-regular graphs and edges. Notice this bound is still tight for all $n$, as demonstrated by paths. Here if $G$ is $H$-free, then $G$ does not have $H$ as an induced subgraph. Meanwhile, as is standard, the claw denotes the graph $K_{1,3}$.
@@
The final set of relaxations of Conjecture 1.2 we consider is to study covers by cycles and edges, rather than decompositions into cycles and edges. Recall that Pyber [24] previous proved that $\operatorname{g_{ce}}(G)\leq n-1$ for all graphs. Note this upper bound is still best possible due to trees. We are able to show the only graphs to require n1n-1 cycles or edges in a cover are trees.
@@ 6. Future Work
Show that if $G$ has maximum degree $5$, or if $G$ is $6$-regular, or $8$-regular, then $G$ can be decomposed into at most n1n-1 cycles and edges.
@@
For all $n$ vertex graphs $G$, $\operatorname{f_{re}}(G)=O(n)$.
@@ Acknowledgement
The research visit of S. Akbari at Simon Fraser University was supported in part by the ERC Synergy grant (European Union, ERC, KARST, project number 101071836). Clow is supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) through PGS D-601066-2025. We would also like to thank Bertille Granet and Matija Bucić for independently making us aware of Erdős and Gallai’s conjecture, as well as the literature regarding it, following the released of the first version of this paper.